### v2

# AbbIE: Autoregressive Block-Based Iterative Encoder for Efficient Sequence Modeling

Abstract
We introduce the Autoregressive Block-Based Iterative Encoder (AbbIE), a novel recursive generalization of the encoder-only Transformer architecture, which achieves better perplexity than a standard Transformer and allows for the dynamic scaling of compute resources at test time. This simple, recursive approach is a complement to scaling large language model (LLM) performance through parameter and token counts. AbbIE performs its iterations in latent space, but unlike latent reasoning models, does not require a specialized dataset or training protocol. We show that AbbIE upward generalizes (ability to generalize to arbitrary iteration lengths) at test time by only using 2 iterations during train time, far outperforming alternative iterative methods. AbbIE’s ability to scale its computational expenditure based on the complexity of the task gives it an up to 12% improvement in zero-shot in-context learning tasks versus other iterative and standard methods and up to 5% improvement in language perplexity. The results from this study open a new avenue to Transformer performance scaling. We perform all of our evaluations on model sizes up to 350M parameters.
1 Introduction
Transformer architectures have become foundational in modern machine learning, enabling state-of-the-art performance across diverse domains (Radford et al., 2019; Jumper et al., 2021). The Transformer’s performance has been shown to depend on the size of the network (parameter count) (Hoffmann et al., 2022; Kaplan et al., 2020; DeepSeek-AI et al., 2024), the amount of data used for training (Hägele et al., 2024; Sardana et al., 2025), and to a degree on the quality of the data the network is trained on (Allal et al., 2025a; Penedo et al., 2024a).
Scaling performance remains challenging due to hardware limitations. In particular, the slower growth of GPU memory capacity compared to computational throughput presents a significant bottleneck to increasing Transformer size for improved performance (Wolters et al., 2024). Different methods have been developed to alleviate this challenge. Examples include quantizing models into lower precisions (Wang et al., 2024), distilling larger models into smaller ones (Hsieh et al., 2023), and pruning parts of the network (Lu et al., 2024). Another popular way to circumvent the memory constraints of Transformers is by using a Mixture of Experts (MoE) architecture (Fedus et al., 2022), which decreases memory requirements by activating only certain parts of the model during inference (Pope et al., 2022; Du et al., 2022).
Figure 1: Performance comparison of AbbIE-D (our method) versus baseline models STD and 0ptacross four benchmarks (Hellaswag, Lambada, ARC-Easy, and CommonsenseQA) for 350M model. Results demonstrate consistent accuracy improvements with increasing training iterations (1, 2, 4, and 8), showing our model significantly outperforms baselines across all benchmarks.
Recently, recurrent formulations of Transformers have appeared as a promising candidate to continue scaling performance without increasing memory (Geiping et al., 2025; Yang et al., 2024). These formulations have shown that the Transformer can operate several times on its token stream, iteratively, to improve performance (Yang et al., 2024). Iteration is employed during both training and inference to allow scaling of resources at test time. This mode of operation is distinctly different from other seemingly recursive language models such as selective state space models (Mamba) (Gu and Dao, 2023) or recursive formulations of the attention mechanism (RKWV) (Peng et al., ). Recurrent Transformers have also been shown to have superior reasoning capabilities, sometimes outperforming models with an order of magnitude more parameters (Yang et al., 2024). However, recurrent Transformers usually rely on a large number of iterations to generalize during training and are often bound to a particular task (Bae et al., 2024; Giannou et al., 2023; Yang et al., 2024). As a result, they are not suitable as drop-in replacements for standard Transformer architectures. The high iteration count also leads to reduced throughput (as the total number of parameters being inferred becomes much larger than the number of unique parameters) and increased training cost. To address these limitations, we propose the Autoregressive block-based Iterative Encoder (AbbIE), an iterative formulation of the encoder-only Transformer. AbbIE leverages the structure of a Transformer’s latent space achieve better perplexities and downstream task performance, but unlike latent reasoning models (Coconut) (Hao et al., 2024) does not require a specialized dataset. Two variants emerge, namely AbbIE-C and AbbIE-D. The difference between the two is in way they handle input injection (Fig. 3). We study and compare both variants. Unlike previous iterative architectures, AbbIE is a generalization of the Transformer network. It can be used as a direct replacement: it matches the performance of a standard Transformer after a single iteration, and requires only two iterations during training to enable lower perplexity and higher ICL scores (Fig. 1) when scaling iterations at test time. This allows AbbIE to be a viable alternative to standard Transformers even when scaling test time compute is not necessary or possible.
Through extensive experiments, we show that AbbIE consistently outperforms standard Transformers for a fixed token budget. When targeting a fixed perplexity at token amounts beyond the compute-optimal point, AbbIE approaches the FLOP cost of a standard Transformer, whereas previous recursive methods have a significantly higher FLOP cost. This makes AbbIE a computationally efficient addition to large-scale training runs, as typically modern Transformers train to well above $20\times$ the compute-optimal token count (Allal et al., 2025b; DeepSeek-AI et al., 2024). We demonstrate that by only training on two iterations, we enable AbbIE to scale to arbitrary iteration numbers with lower perplexity and improve performance on a range of in-context learning (ICL) tasks during inference. This property allows AbbIE to match the performance of a standard Transformer, while retaining the flexibility to scale inference compute when beneficial. To our knowledge, AbbIE is the only recurrent Transformer which possesses these characteristics. They make our method readily adoptable, as it functions equivalently to a standard Transformer, unless test-time scaling is explicitly desired. In summary, the contributions of this paper are:
  • A new recurrent method for auto-regressive Transformers, termed AbbIE, that can scale performance during test time.
  • A training methodology for recurrent Transformers networks, which requires only two iteration at train time.
  • Extensive experiments on language modeling tasks that show AbbIE outperforms standard Transformers for the same token budget.
  • Extensive experiments on zero-shot In-Context Learning (ICL) tasks that show AbbIE can improved the performance past the number of training iterations.
2 Autoregressive block-based Iterative Encoder: AbbIE
In this section, we present the architecture and operational mechanism of AbbIE.
2.1 Design
Figure 2: Architecture of AbbIE method. The model consists of three main components: Head (green), Body (blue), and Tail (yellow). Input embeddings first pass through the Head, where they are converted from token space to concept space. The embeddings are then iterated over $r$ times (k=0,1,,r1)(k=0,1,...,r-1) by the Body. The final embeddings are then processed by the Tail and projected back into token space. $N_{h}$$N_{b}$ and $N_{t}$ show the number of Transformers blocks in each component.
AbbIE consists of three structural groups and follows the structure of a decoder-only architecture (Vaswani et al., 2017; DeepSeek-AI et al., 2024). We term the three groups Head, Body, and Tail. This separation is similar to the structure presented by Geiping et al. (2025), but unique in that AbbIE does not require additional projection matrices or large training iterations. A visual representation of the separation is shown in Fig. 2. The primary distinction between AbbIE and a standard Transformer lies in the iterative structure of the Body: instead of processing the tokens once, AbbIE applies the body stack repeatedly across iterations. In AbbIE, each structural group contains $N_{h}$, $N_{b}$, $N_{t}$ Transformer blocks respectively. The Body further operates recursively over $r$ iterations. This structural grouping can be seen as a generalization of the original Transformer architecture, which emerges as a special case under our formulation when the iteration number $r=1$. The Head contains the embedding layer $W$ and $N_{h}$ Transformer blocks. Its role is to transforms the embedded token representations into concept space through the detokenization process (Kaplan et al., 2024). Concept space, is an internal space in the Transformer body, described by Kaplan et al. (2024); Hao et al. (2024). Once in concept space, the Body iterates over the residual stream using one of two operational variants: AbbIE-C or AbbIE-D. In AbbIE-C, during each iteration, the Body operates directly on the output of the previous iteration. In AbbIE-D, the outputs of each iteration are added to the original input, forming a residual chain. Finally, the Tail maps the representations back to token space before unembedding.
2.2 Operation
We adopt a recurrent formulation to allow the model to spend a variable amount of computation per token. Prior work has demonstrated that allowing Transformers to reprocess their residual streams across multiple passes improves performance (Nye et al., 2021; Hao et al., 2024). AbbIE has two operational variants: a continuous chain-of-thought inspired mode (AbbIE-C) and a diffusion-inspired mode (AbbIE-D). The methods differ in the way they use the residual stream between iterations in the Body. We motivate the creation of both methods through the Path Independence (PI) property (Anil et al., 2022). PI, introduced by Anil et al. (2022), is a relaxed form of a contraction mapping (lacks strictly decreasing requirement). Formally, a function $f$ is PI if repeated application it to an input $x$ leads to a fixed point $z^{*}$:
$f^{(\infty)}(x,z_{0})=z^{*},\text{for any}\leavevmode\nobreak\ z_{0}$ (1)
where $z_{0}$ is some initial state and the superscript (⋅) indicates repeated application. PI has been shown to correlate with a model’s ability to generalize across iteration counts, i.e. upwards generalization. While Anil et al. (2022) identify the necessary components for PI (dependence on the original input at each iteration and the injection of noise in the first iteration), these do not preclude other methods from achieving fixed points and thus upwards generalization. In this work, we omit $z_{0}$ or implicitly set $z_{0}=\mathbf{0}$. Therefore we can use a formulation which only relies on the input as:
$f^{(n)}(x_{n})\approx f^{(n+1)}(x_{n+1})\leavevmode\nobreak\ \text{as}\leavevmode\nobreak\ n\to\infty$ (2)
where the subscript shows the iterant index. Therefore, we are only concerned with the inclusion of inputs into each iteration. Transformers inherently depend on the original inputs via the residual stream, suggesting that simply looping within concept space may suffice for generalization. Based on this idea, we define our first variant, AbbIE-C, as direct recursion over the residual stream (Fig. 3), akin to continuous chain-of-thought models (Hao et al., 2024). Formally, we can define AbbIE-C as:
$h_{k+1}=B(h_{k}),\quad\text{for }k=0,\ldots,N-1$ (3)
where $B(\cdot)$ is the Body component. The second variant, AbbIE-D, re-injects inputs into each block through an inter-iteration residual stream. Formally, this can be defined as:
$h_{k+1}=B(h_{k})+h_{k},\quad\text{for }k=0,\ldots,N-1$ (4)
Since each of the Transformer blocks in the Body has a residual connection, AbbIE-C and AbbIE-D could seem equivalent. The additional residual connection around the whole Body block, however, reinforces the relative strength of the information contained in the input embeddings $h_{0}$ compared to the cumulative updates originating from the Attention and FFN components in the Transformer blocks. To see this, we can define the delta between iteration states induced by $B(\cdot)$ as:
$\Delta h_{k}=\sum_{i=0}^{N_{b}-1}(\text{Attn}_{i}+\text{FFN}_{i})$ (5)
where each element in the sum depends on all the previous elements. Then, for AbbIE-C, from (3), the subsequent iteration output becomes $h_{k+1}=h_{k}+\Delta h_{k}$ while for AbbIE-D, from (4), it is $h_{k+1}=2h_{k}+\Delta h_{k}$. To achieve the same output composition, the model should learn to make $h_{k}[\texttt{AbbIE-D}]=h_{k}[\texttt{AbbIE-C}]/2$, but then this would make $\Delta h_{k}[\texttt{AbbIE-D}]$ different from $\Delta h_{k}[\texttt{AbbIE-C}]$ in an unpredictable, non-linear way.
Figure 3: Architectural comparison of the two AbbIE variants. AbbIE-C (left) uses the Transformer’s existing residual connection as means of input injection, whereas AbbIE-D (right) introduces and additional inter-iteration residual. Although similar it is not possible, given an arbitrary AbbIE-D instance, to produce an AbbIE-C instance that computes the same function.
3 Experimental Setup
In this section we describe the training procedures used throughout this work, along with the baselines and evaluations used to assess model performance.
3.1 Baselines
To evaluate the effectiveness of AbbIE, we conduct comparisons against both standard and recursive Transformer baselines. 0ptis our implementation of the recursive method proposed by Geiping et al. (2025). Like AbbIE, 0ptapplies iteration over a shared set of Transformer blocks in latent space. However, instead of relying on the residual stream for input re-injection, 0ptconcatenates the input embeddings to, initially a randomly sampled vector, and in subsequent iterations to the output of the previous iteration. This concatenated vector is projected back to the model dimension. While the original method uses non-standard width-to-depth ratios, we use GPT-2-like architectural proportions to ensure fair comparison. Although Geiping et al. (2025) propose a custom sandwich normalization strategy, they report that at small scales, most normalization approaches (pre-norm, post-norm, etc.) perform comparably. Accordingly, we adopt a consistent pre-norm configuration across all baselines.
For our standard Transformer baseline (Std), we use a decoder-only architecture loosely inspired by the design choices in Grattafiori et al. (2024), with proportions similar to those used in GPT-2 (Radford et al., 2019). This model lacks any form of iteration and serves as the canonical single-pass baseline.
3.2 Training Configuration
All models, defined in Table 1, are trained using standard Transformer training practices. We define the compute-optimal token budget (COT) as 20 tokens per model parameter, consistent with established empirical scaling laws (Hoffmann et al., 2022). All models are trained up to the COT budget, unless specified otherwise. In specific analyses involving perplexity scaling, we include training runs beyond the COT budget, but all baseline comparisons are made at the compute-optimal point. Training is performed using the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We use two global batch sizes: 256 samples for the small-batch (SB) regime and 1024 samples for the large-batch (LB) regime. The two regimes allows us to test how recurrent processing interacts with optimization dynamics under different gradient noise conditions McCandlish et al. (2018). All models are trained on fixed-length sequences of 2048 tokens. We adopt a Warmup-Stable-Decay learning rate schedule. We use linear warmup lasting $0.25\times\text{COT}$ steps. We use a cosine decay over $0.20\times\text{COT}$ steps, in line with best practices from DeepSeek-AI et al. (2024). Tied input-output embedding matrices are used across all models, ensuring that observed improvements stem from the model body itself rather than bi-gram memorization or shortcut learning. Architecturally, we adopt a modern Transformer design with Grouped Query Attention (GQA) (Ainslie et al., 2023; Grattafiori et al., 2024), Rotary Positional Embeddings (RoPE) (Su et al., 2024), SiLU activation (Elfwing et al., 2017), Pre-norm residual structure (Xiong et al., 2020) and RMSNorm(Zhang and Sennrich, 2019).
Table 1: Model configurations for evaluations across two model sizes and two batch sizes. We use GPT-2-like architectures. $N_{h,b,t}$ show component numbers, $d_{model}$ and FFN size define the width of the network. Attn heads and KV heads define the GQA paramters and RoPE $\theta$ shows the base used in the rotary positional embeddings.
Size BS $N_{h}$ $N_{b}$ $N_{t}$ $d_{model}$ FFN size Attn heads KV heads RoPE $\theta$
200M 256 2 10 2 1024 4096 16 8 10,000
200M 1024 2 10 2 1024 4096 16 8 10,000
350M 1024 2 20 2 1024 4096 16 8 10,000
For the smaller size models we have performed 5 training runs each with different seeds to verify the statistical significance of our experimentation; however for the large models we have only conducted a single training run. Best performance is reported. No hyperparameter tuning is performed per variant. All experiments were run on a single H100 GPU.
3.3 Training Dataset and Evaluation Metrics
We train on a mixture of datasets proposed by von Platen and Wolf (2023) which provides models with a high-quality mix to achieve a high in-context learning performance. We use 80% of FineWeb-Edu Penedo et al. (2024b), 10% Cosmopedia-v2 Ben Allal et al. (2024) and 10% Python-Edu Kocetkov et al. (2022). FineWeb-Edu is a filtered subset of the broader FineWeb dataset, curated to emphasize high-quality, diverse, and educationally valuable text. Cosmopedia-v2 consists of synthetic encyclopedic content generated to enhance factual recall and structured knowledge. Python-Edu contains educational programming content, with a focus on natural language explanations of Python code and logic.
For evaluation, we focus on a suite of in-context learning (ICL) benchmarks particularly in the domains of commonsense reasoning and language understanding. These benchmarks are: HellaSwag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), ARC-Easy (Clark et al., 2018) and CommonsenseQA (Talmor et al., 2019). At the model sizes that we examine, some ICL tasks are unsuitable for evaluation as they do not improve above the random baseline level. One of these tasks is CommonsenseQA which we intentionally include to show that iteration scaling significantly improves our reasoning capabilities. Each evaluation is performed in the zero-shot setting. We report accuracy as the primary metric across all tasks. Our goal in selecting these benchmarks is to assess effectiveness of iterative refinement mechanisms in reasoning-centric tasks that benefit from multi-step internal processing Biran et al. (2024).
4 Results
In this section we present the results from evaluating the AbbIE method, comparing it against current approaches in recurrent and standard Transformers. With these experiments we answer the following research questions:
RQ1: Can AbbIE achieve a fixed point iteration by using the residual as input injection? (Sec. 4.1)
RQ2: Can AbbIE reach a lower perplexity for the same size and token budget? (Sec. 4.2)
RQ3: Can AbbIE achieve upward generalization with only two training iterations? (Sec. 4.3)
RQ4: Can AbbIE’s scale test time performance? (Sec. 4.3)
4.1 Fixed Point (RQ1)
To evaluate whether AbbIE converges to a fixed point in iteration count, we measure the distance between the input and output embeddings at each iteration. This analysis is performed on randomly sampled data points from the ICL evaluation sets. As a control, we include the 0pt method, which is known to be path independent by construction and thus guaranteed to reach a fixed point. Figure 4 summarizes the results. We observe that AbbIE-D exhibits smooth decay past a certain iteration, eventually converging. This behavior is shared with the 0ptmethod showing that AbbIE-D, in the samples we have evaluated on, reaches a fixed point. In contrast, AbbIE-C fails to converge after 32 iterations. This shows that the residual steam, present in the body of Transformer, is in itself not sufficient and an extra inter-iteration residual connection is needed. Due to AbbIE-C inability to converge it is excluded from further analyses.
Figure 4: Inter-iteration distances vs number of iterations for 16 random samples from various benchmarks (Hellaswag, Lambada, ARC-Easy, and CommonsenseQA). We see that AbbIE-D (top) and 0pt(bottom) reach a fixed point. This implies greater stability and better upward generalization capabilities. AbbIE-C (middle) consistently diverges.
4.2 Performance scaling and computational expenditure (RQ2)
To evaluate the performance of AbbIE we show how perplexity evolves as a function of token budget. Figure 5(b) summarizes this comparison. Across all configurations, AbbIE consistently outperforms both the standard Transformer and 0ptbaseline in terms of perplexity.
AbbIE follows scaling trends consistent with standard Transformers, showing steady reductions in perplexity with increased model size and token count. This indicates that AbbIE retains the favorable scaling properties of Transformer architectures while offering improved perplexity.
Next, we evaluate the relationship between computational cost and model performance during training. As other recurrent Transformers require many iterations to train, we find it important to demonstrate that the training costs of AbbIE are much smaller than a typical recurrent formulation and are comparable with standard Transformers. Following the methodology from Hoffmann et al. (2022), we estimate the floating point operations (FLOPs) required to reach various perplexity thresholds. Results are shown in Figure 5(a). AbbIE follows the same scaling laws as standard Transformers: lower perplexities require more computation. However, it generally incurs a higher FLOP cost. Despite this, the additional cost decreases over time. To show this we define FLOP efficiency as the ratio of AbbIE FLOPs to standard Transformer FLOPs. The gradient of FLOP efficiency in Figure 5(a) is negative. Therefore, for very long training runs, on the order of $20\times$ to $200\times$ the compute-optimal budget, which is common in large-scale training, we hypothesize the total computational cost of AbbIE will converge to that of the standard Transformer. This is because AbbIE maintains a consistent perplexity advantage, while the marginal improvements in perplexity become harder to achieve as training progresses. Crucially we show that AbbIE-D’s improvements in perplexity and scaling stability are valid across model and batch sizes.
4.3 Test time scaling (RQ3, RQ4)
(a) Computational expenditure of AbbIE and its scaling with perplexity (Top). FLOP efficiency (Bottom), defined as the ratio of FLOPs spent on AbbIE to FLOPs spent on Std, has a negative gradient. This shows that with long training runs, the FLOP efficiency will improve.
(b) Perplexity vs token count follows standard scaling law patterns, with AbbIE having a roughly 5% better perplexity than Std. 0ptfollows the same trend, but with a worse perplexity.
Figure 5: Performance scaling and computational expenditure of AbbIE vs Std vs 0pt. The figures shows that AbbIE follows typical scaling laws and has an improved perplexity. For token amounts near COT, the computational cost of AbbIE is typically higher than that of Std but lower than 0pt.
(a) Comparison of perplexity across iterations for AbbIE-D models of different sizes. The 350M parameter model continues to improve perplexity past the number of trained iterations. This shows that it has generalized in iteration count.
(b) Comparison of perplexity (log scale) across iterations for AbbIE-D, 0ptand Std. AbbIE-D exhibits stable perplexity, whereas 0ptexperiences perplexity collapse at $r\neq 2$. Std is included to show the behavior of a non-iterative model in a iterative setting.
Figure 6: Perplexity trends across iteration counts for different model variants. Sub figure (a) shows a detailed comparisons between AbbIE-D configurations.
To evaluate the scaling characteristics of AbbIE-D we measure the perplexity at different iteration counts. We compare versus the 0ptbaseline. Figure 6(a) shows that AbbIE-D possesses a stable perplexity in the number of iterations. The 200M parameter versions achieve their lowest perplexity at $r=2$, which matches their training regime. However, the 350M parameter version achieves its lowest perplexity at $r=4$, twice the training regime. This suggests that AbbIE-D at the larger size was able to successfully upward generalize (RQ3). This finding is also confirmed by the ICL task performance. In contrast, Figure 6(b) shows that 0ptcannot generalize in the same training regime.
Figure 1 shows the ICL task scaling of AbbIE-D compared to Std and 0pt(only at $r=2$, where perplexity is stable). We show that AbbIE-D’s perplexity scaling has translated into improved ICL task performance. Even when the perplexity at $r=8$ is higher than the minimum the ICL task performance continues to improve. This is important, as most recurrent Transformers’s performance plateaus at the number of training iterations. Conversely, we see that once generalization is achieved in AbbIE-D, even though perplexity begins to increase after a certain point, ICL task performance continues to improve. This also shows that the relationship between ICL and perplexity is not strict. Notably we show that AbbIE-D continues to improve performance at $4\times$ the training iteration count, which no other general recursive transformer has shown. (RQ4)
For AbbIE-D 350M we see improvements of up to 12% on HellaSwag. In CommonsenseQA at 8 iterations we see a performance significantly higher than the random baseline which was achieved by all other models. However, as no other method was able to surpass the random baseline, we cannot claim direct percentage improvement in CommonsenseQA. However, we show that AbbIE improves the emergent reasoning capabilities of the Transformer. The magnitude of the ICL task performance is on par with other models of a similar size trained to COT AMD (2023); von Platen and Wolf (2023). The rest of the comparisons are summarized in Table 2. For AbbIE-D 200M we closely match the performance of the standard Transformer with a distinctive reduction of performance at $r=1$. However, this difference disappears for AbbIE-D 350M, which suggest some critical model size larger than 200M has to be used for the model to have the spare capacity to learn to improve the residual stream with iterations. The performance gains also remain across different batch sizes and therefore gradient noise regimes. This indicates that the AbbIE approach is robust to the sharp minima present in larger batch training.
Table 2: Zero-shot performance (%) across recursive iterations on standard In-Context Learning (ICL) benchmarks. We report the best result from each method-category. AbbIE-D consistently outperforms all other approaches and is the only method to achieve a CSQA score (For the Commensense Question Answering task) above the random baseline. From the results, we can see 0ptachieves $29.6$ for 200M model. But by comparing it against Fig. 6(b) we can observe that the model’s perplexity has collapsed (indicating model collapse). Therefore, the accuracy of $29.6$ potentially is a random outcome.
Size BS Model HellaSwag LAMBADA ARC-E CSQA
200M 256 Std 30.1 24.2 45.6 20.0
AbbIE-D ($r=1$) 30.2 6.0 42.5 20.0
AbbIE-D ($r=2$) 31.4 23.1 45.9 20.0
AbbIE-D ($r=4$) 31.3 24.2 48.7 20.0
AbbIE-D ($r=8$) 31.3 22.7 49.2 20.0
0pt($r=1$) 25.0 0.0 26.8 20.0
0pt($r=2$) 29.7 22.2 46.3 20.0
0pt($r=4$) 25.0 0.0 27.5 29.6
0pt($r=8$) 25.0 0.0 25.0 20.0
1024 Std 30.2 23.2 47.0 20.0
AbbIE-D ($r=1$) 29.2 11.0 41.8 20.0
AbbIE-D ($r=2$) 30.0 21.8 45.6 20.0
AbbIE-D ($r=4$) 30.4 25.1 46.5 20.0
AbbIE-D ($r=8$) 30.6 25.9 47.3 20.0
0pt($r=1$) 25.0 0.0 26.5 20.0
0pt($r=2$) 29.4 22.5 45.6 20.0
0pt($r=4$) 25.0 0.0 28.5 29.6
0pt($r=8$) 25.0 0.0 25.8 20.0
350M 1024 Std 32.5 26.0 48.2 20.0
AbbIE-D ($r=1$) 31.9 25.0 48.5 20.0
AbbIE-D ($r=2$) 33.6 27.8 49.0 21.0
AbbIE-D ($r=4$) 34.5 28.0 50.8 25.0
AbbIE-D ($r=8$) 34.8 28.0 51.3 27.6
0pt($r=1$) 25.0 0.0 25.4 20.0
0pt($r=2$) 32.5 26.5 49.3 20.0
0pt($r=4$) 25.0 0.0 28.5 20.0
0pt($r=8$) 25.0 0.0 26.4 20.0
Random Baseline 25.0 0.0 25.0 20.0
5 Background and Related Work
Latent space Transformers perform the majority of their computation in a high-dimensional latent space. While the architecture of a standard Transformer is uniform across layers, the representations it produces are not. Recent work has shown that Transformer latent space can be broadly divided into two distinct regions Kaplan et al. (2024); Hao et al. (2024). The first, which we refer to as token space consists of the low-level hidden states produced by the embedding matrix, where each vector corresponds to a subword unit generated by the tokenizer. The second, which we refer to as the concept space, consists of higher-level semantic representations, which evolve through the depth of the Transformer Katz and Belinkov (2023). These are ultimately mapped back into token space by the final Transformer blocks and the output (enembedding) layer (Kaplan et al., 2024). The transition from token space to concept space, referred to as detokenization (Kaplan et al., 2024; Elhage et al., 2022; Gurnee et al., 2023) occurs within the early layers of the model. Within concept space, several studies have demonstrated that Transformer blocks can be reordered, skipped, or reused (Fan et al., 2019; Elhoushi et al., 2024; Shen et al., 2025) without significantly impacting performance. This suggests that individual layers specialize in specific conceptual transformations, rather than relying on rigid positional dependencies.
Path independence Path independence, introduced by Anil et al. (2022), is a property of an iterative model to converge to a fixed point regardless of its initialization. The authors argue that PI strongly correlates with high accuracy on out-of-distribution samples (upward generalization). Intuitively, convergence to a fixed point reflects stable refinement of representations; in contrast, non-convergent models may exhibit erratic updates that harm performance. Formally, a method is path independent if it converges to the same limiting behavior from any initial state Anil et al. (2022). In this work, we use concepts from path independence (input injection) and empirically show the existence of a fixed point by analyzing the sequence of representations across iterations.
Recurrent Transformers Recurrent formulations of Transformer networks have gained attention as a promising approach to overcome the scaling limitations of standard architectures. They benefit from increased computation depth and improve performance on certain tasks Geiping et al. (2025). These methods typically use weight tying across layers; some, such as the Universal Transformer (Dehghani et al., 2019), also include depth embeddings, which can limit scalability at test time. Others incorporate inter-iteration projections to sustain performance at increasing depths. A common approaches in recurrent Transformers is the re-injection of inputs at each iteration, often coupled with random noise (Geiping et al., 2025; Yang et al., 2024; Giannou et al., 2023) to aid in the performance and extrapolation capabilities of the model. The method proposed by Geiping et al. (2025) shares architectural similarities with AbbIE but differs in several key aspects. Their model initializes the latent space with random vectors and is both trained over a range of randomly sampled iteration count. As a result, its performance saturates near the average number of training iterations. In contrasts, AbbIE leverages the internal structure of the Transformer’s latent space and residual connections to generalize across arbitrary iteration counts without the need for specialized embeddings or projections. Additionally, AbbIE is trained with just two iterations and achieves stable perplexity at test time regardless of depth, offering a more efficient and generalizable approach.
6 Discussion and Limitations
Recurrent Transformer architectures have demonstrated strong performance on a range of downstream tasks, outperforming non-recurrent models. However, these methods are either highly modified versions of the Transformer network, which are unsuitable for general language modeling or have not been evaluated in a controlled, one-to-one comparison against identical standard Transformer architectures. As a result, it remains unclear how much of the observed gains are due to recurrence itself versus other architectural or training differences. A major limitation of existing recurrent models is their consistently larger computational cost. They require multiple iterations per token and lack the ability to adaptively reduce computation when recurrence is unnecessary. This restricts their applicability in resource-constrained settings and reduces throughput. In this work, we introduced a new approach, AbbIE, that achieves upwards generalization with as few as two iterations. Although lower than all other recurrent Transformers, for token budgets close to COT, AbbIE still uses more FLOPs. Our method produces lower perplexity with increasing loop count, enabling dynamic control over compute at inference time. Our experimental results show that AbbIE reaches a 5% improvement in perplexity and a upward of 12% improvement in ICL tasks with improved emergent properties. When run with a single iteration, our model matches the perplexity of a standard Transformer. We demonstrate that the representations in our method gradually converge toward a fixed point. However, prior work suggests that Transformer dynamics may not align perfectly with fixed-point behavior. This warrants further investigation into the nature of iterative refinement in attention-based models. Future work will explore enhanced training objectives and normalization strategies to further improve perplexity and convergence.